A 1.0- L buffer solution contains 0.100 molH C 2 H 3 O 2 and 0.100 molNa C 2 H 3 O 2 . The value of K a for H C 2 H 3 O 2 is 1.8× 10 −5 .

Part A

Calculate the pH of the solution upon addition of 20.9 mL of 1.00 MHCl to the original buffer.
Answer #1

mol of HCl added = 1.0M *0.0209 L = 0.0209 mol

C2H3O2- will react with H+ to form HC2H3O2

Before Reaction:

mol of C2H3O2- = 0.1 mol

mol of HC2H3O2 = 0.1 mol

after reaction,

mol of C2H3O2- = mol present initially - mol added

mol of C2H3O2- = (0.1 - 0.0209) mol

mol of C2H3O2- = 0.0791 mol

mol of HC2H3O2 = mol present initially + mol added

mol of HC2H3O2 = (0.1 + 0.0209) mol

mol of HC2H3O2 = 0.1209 mol

Ka = 1.8*10^-5

pKa = - log (Ka)

= - log(1.8*10^-5)

= 4.745

since volume is both in numerator and denominator, we can use mol instead of concentration

use:

pH = pKa + log {[conjugate base]/[acid]}

= 4.745+ log {7.91*10^-2/0.1209}

= 4.56

Answer: 4.56